# Online Learning for Supervisory Switching Control

### 16 Mar 2026 v1

@@ Abstract
We study supervisory switching control for partially-observed linear dynamical systems. The objective is to identify and deploy the best controller for the unknown system by periodically selecting among a collection of $N$ candidate controllers, some of which may destabilize the underlying system. While classical estimator-based supervisory control guarantees asymptotic stability, it lacks quantitative finite-time performance bounds. Conversely, current non-asymptotic methods in both online learning and system identification require restrictive assumptions that are incompatible in a control setting, such as system stability, which preclude testing potentially unstable controllers. To bridge this gap, we propose a novel, non-asymptotic analysis of supervisory control that adapts multi-armed bandit algorithms to address these control-theoretic challenges. Our data-driven algorithm evaluates candidate controllers via scoring criteria that leverage system observability to isolate the effects of historical states, enabling both detection of destabilizing controllers and accurate system identification. We present two algorithmic variants with dimension-free, finite-time guarantees, where each identifies the most suitable controller in $\mathcal{O}(N\log N)$ steps, while simultaneously achieving finite $L_{2}$-gain with respect to system disturbances.
00footnotetext: This work was supported by the ONR grant #N00014-23-1-2299.00footnotetext: Corresponding author: Haoyuan Sun (haoyuans [at] mit [dot] edu).
@@ 1 Introduction
This paper presents a supervisory switching framework that guarantees finite-time adaptation for unknown linear systems, leveraging online algorithm to establish non-asymptotic analysis for this classical problem. In many applications involving complex systems, such as power systems [16], autonomous vehicles [2], and public health [5], it is impractical for a single controller design to achieve acceptable performance across all operating conditions of the unknown plant. To address this challenge, switching control employs a collection of candidate controllers alongside a switching policy that selects the controller best suited to the current system behavior [14]. This approach has important applications in many established control paradigms, including gain scheduling [25, 23] and linear parameter-varying control [32, 17].
@@
To develop a non-asymptotic analysis of supervisory switching control, we revisit this classical problem through the lens of modern learning techniques. A natural starting point is the non-asymptotic analysis of system identification, which estimates the parameters of an unknown dynamical system from a single trajectory of input-output data. Conceptually, estimator-based supervisors select the model that minimizes the error between predicted and observed outputs, akin to the least-squares objectives underpinning system identification. Recent advances in non-asymptotic system identification have extended classical asymptotic results [15] to quantify the sample complexity of learning accurate system estimates from data [28, 33, 13]. This has led to numerous methods for the non-asymptotic identification of partially observed linear systems, typically by estimating the system’s response matrix to an exploratory signal and reconstructing the system parameters via the Ho-Kalman algorithm [21, 26, 24, 4]. However, these methods cannot be directly applied to our supervisory switching control paradigm because they presuppose closed-loop stability. This assumption is frequently violated in switching systems, where not all candidate controllers stabilize the underlying plant. Therefore, standard system identification approach cannot ensure safe exploration, as the potentially unstable controllers may cause the system to diverge before sufficient data is collected.
@@
In this work, we propose a data-driven analysis of supervisory switching control that establishes non-asymptotic guarantees for adapting to partially-observed linear systems with a collection of $N$ candidate controllers. Drawing on techniques from the machine learning and multi-armed bandit literature, our approach (Algorithm 1) evaluates each candidate controller over fixed switching intervals. Specifically, our evaluation criteria leverage system observability to approximate state trajectories from output data, effectively isolating the influencing of earlier destabilizing controls. This eliminates the need for the closed-loop stability or fully-observed state assumptions mandated by previous works.
@@ 3 Algorithm Design and Evaluation Criteria
2:Define $\bar{s_{i}}(q)=\frac{1}{q}\sum_{\ell=1}^{q}s_{i}[\ell]$ be the average score
3:Let $Q_{i}(\ell)$ be the number of applications of controller $i$ by episode $\ell$
@@
  • Theoretical analysis of online learning with states — such as reinforcement learning (RL) — typically rely on one of two assumptions to bound state trajectories during exploration. Episodic RL [9] assumes the system is periodically reset to a default initial state distribution, preventing compounding errors across episodes. Conversely, non-episodic online RL analyzes continuous operation without resets [27, 20], but relies on restrictive structural assumptions, such as open-loop system stability, to prevent the state from diverging. In contrast, our setting requires exploring potentially destabilizing controllers without resets. This means that the systems states may diverge, thus violating the core assumptions of standard RL analysis.
@@ 3.1 Criterion 1: Instability Detection
For convenience, denote the current episode’s outputs as $y_{1},\dots,y_{\tau}$ and let $j$ index the active controller. Under Assumption 1, the closed-loop system is $\nu$-strictly observable, which yields the following input-output relation:
@@
We first consider the case where the active controller matches the true system ($i_{\star}=j$). In this case, the closed-loop dynamics are stable, and the Toeplitz matrices are bounded in norm by $\left\lVert C_{[i_{\star},i_{\star}]},A_{[i_{\star},i_{\star}]},B_{[i_{\star},i_{\star}]}\right\rVert_{\mathcal{H}_{\infty}}$ (see [30]). Furthermore, Assumption 1 guarantees $\left\lVert\mathcal{O}_{[i_{\star},i_{\star}]}^{\dagger}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon_{c}$. Consequently, the aggregated disturbance term $\mathcal{O}_{[i_{\star},j]}^{\dagger}\xi$ is sub-Gaussian. By standard concentration inequalities for sub-Gaussian random variables, the estimate
@@
Next, when the active controller is destabilizing, Assumption 2 ensures the closed-loop system possesses an explosive mode in the direction of the leading eigenvector of $A_{[i_{\star},j]}$. Since the system is persistently excited by Gaussian process noise, this explosive mode is activated with non-zero probability. In this regime, the residual $y_{t}-\hat{y}_{t}$ can be decomposed as the sum of estimation error plus accumulated disturbances, where the first part arises because $\hat{x}_{1}$ does not approximate the true initial state $x_{1}$ when $i_{\star}\neq j$. Critically, the estimation error is determined entirely by the first $\nu$ observations, while the system disturbances are accumulated at every step of the trajectory. As a result, the accumulated disturbance contains a component driven by process noise realized after the $\nu$-step, which is independent of the estimation error. Hence, with non-zero probability, these independent components will not cancel, and contributions from the explosive mode will dominate the residual for sufficiently large $\tau$. Thus, the residual $y_{\tau}-\hat{y}_{\tau}$ will often exceed the threshold $\Theta$ when the active controller is destabilizing.
@@ 3.2 Criterion 2: System Identification
3:Compute a virtual trajectory
@@ 4 Main Result
With the guarantees of the two evaluation criteria established, we now analyze the performance of our supervisory switching control algorithm. To balance exploration and exploitation, we draw inspiration from the Upper Confidence Bound (UCB) algorithm [3]. At each episode, we select the controller by maximizing its empirical average score plus with an exploration bonus. The magnitude of this bonus encourages exploration of controllers with limited interactions, and the parameter $a_{\ell}$ controls the exploration strength.
@@
Let $\mu^{+}$ denote a lower bound on the expected score of the matching controller ($j=i_{\star}$), and let $\mu^{-}$ denote an upper bound on the expected score of any non-matching controller ($j\neq i_{\star}$). Similar to standard UCB analysis, if Algorithm 1 selects a non-matching controller $j\neq i_{\star}$ at episode $\ell$, then at least one of the following three conditions must be met:
@@
  • The exploration bonus is too wide: $\exists j\neq i_{\star}$, s.t. $2\sqrt{\frac{a_{\ell}}{Q_{j}(\ell)}}>\mu^{+}-\mu^{-}$.
@@
The first part of this theorem establishes a sample complexity bound that our algorithm requires $\mathcal{O}(N\log N)$ episodes to identify the matching controller with high probability. This represents a significant improvement over our prior work [29] that required $\mathcal{O}(\exp(N))$ steps.
@@
In contrast, our Theorem 3 bounds the undiscounted case ($\lambda=0$). This regime is not covered by classical analysis, as the traditional framework requires a dwell time that scales inversely with $\lambda$. Furthermore, our result provides quantitative bounds for the constants $C_{0},C_{1}$ in terms of the problem parameters, whereas classical analysis is asymptotic and leaves the explicit values of $\overline{C}_{0}$ and $\overline{C}_{1}$ unspecified.
@@ 5 Conclusion
In this work, we presented a supervisory switching control framework that guarantees finite-time adaptation for unknown partially-observed linear systems with a collection of candidate controller. Our proposed data-driven method achieves a dimension-free sample complexity of $\mathcal{O}(N\log N)$ steps to identify the matching controller while maintaining finite $L_{2}$-gain, even when some of the candidate controllers risk destabilizing the underlying plant.
@@ Appendix A Mathematical Preliminaries
Let $\{D_{k}\}_{k=1}^{\infty}$ be a martingale-difference sequence adapted to a filtration $\{\mathcal{F}_{k}\}_{k=1}^{\infty}$. If $D_{k}\in[0,1]$ almost surely for all $k$, then for any $t>0$,
@@
Finally, we introduce the following technical lemma to arrange inequality of the form $b\geq a\log b$, which shows up when we bound the episode length $\tau$.
@@ Appendix B Proof of Proposition 1
3:Compute a virtual trajectory
@@
To start with the proof, we recall that, under Assumption 1, the closed-loop system is $\nu$-strictly observable, which yields the following input-output relation:
@@
$\hat{x}_{1}=\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}y_{\nu}\\ \vdots\\ y_{2}\\ y_{1}\end{bmatrix}=x_{1}+\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}\mathcal{T}^{(\nu-1)}_{[j,j]}\\ \mathbf{0}_{1\times(\nu-1)}\end{bmatrix}\begin{bmatrix}w_{\nu-1}\\ \vdots\\ w_{2}\\ w_{1}\end{bmatrix}+\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}\mathcal{T}^{(\nu-1)}_{[j,j]}\\ \mathbf{0}_{1\times(\nu-1)}\end{bmatrix}\begin{bmatrix}B_{[j,j]}u_{\nu-1}\\ \vdots\\ B_{[j,j]}u_{2}\\ B_{[j,j]}u_{1}\end{bmatrix}+\begin{bmatrix}\eta_{\nu}\\ \vdots\\ \eta_{2}\\ \eta_{1}\end{bmatrix}$.
Under Assumption 1, we have $\left\lVert\mathcal{O}_{[i_{\star},j]}^{\dagger}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon_{c}$. And since the matching controller is stabilizing, the Toeplitz matrices are bounded by $\left\lVert C_{[j,j]},A_{[j,j]},B_{[j,j]}\right\rVert_{\mathcal{H}_{\infty}}$. It follows that $x_{1}-\hat{x}_{1}$ is
@@
Now, we can let $(u_{i_{\star}},v_{i_{\star}})$ be the leading singular vectors of $C_{[i_{\star},j]}A_{[i_{\star},j]}^{\tau-\nu-2}$ corresponding to the top singular value $\geq\varepsilon_{c}\varepsilon_{a}(1+\varepsilon_{a})^{\tau-2\nu-2}$. Since $w_{\nu+1}$ is Gaussian, we know that from standard normal CDF, $\Pr(|v_{i_{\star}}^{\top}w_{\nu+1}|>4\sigma_{w})>4/5$. Then we have that
@@
this criteria correctly detect an unstable $A_{[i_{\star},j]}$ with probability at least 2/5.
@@
this criteria correctly detect an unstable $A_{[i_{\star},j]}$ with probability at least 2/5.
@@ Appendix C Proof of Proposition 2
Then, this criteria considers the active controller’s associated model as matching the true system if $\Delta_{u_{k},v_{k}}$ is small along every critical direction. And this criteria declares non-matching if one of the $\Delta_{u_{k},v_{k}}$ is large.
@@ Appendix D Proof of Theorem 3
Given Proposition 1 and 2, we can instantiate the following guarantees our evaluation metric:

### 17 Mar 2026 v2

@@ Abstract
We study supervisory switching control for partially-observed linear dynamical systems. The objective is to identify and deploy the best controller for the unknown system by periodically selecting among a collection of $N$ candidate controllers, some of which may destabilize the underlying system. While classical estimator-based supervisory control guarantees asymptotic stability, it lacks quantitative finite-time performance bounds. Conversely, current non-asymptotic methods in both online learning and system identification require restrictive assumptions that are incompatible in a control setting, such as system stability, which preclude testing potentially unstable controllers. To bridge this gap, we propose a novel, non-asymptotic analysis of supervisory control that adapts multi-armed bandit algorithms to a control-theoretic setting. The proposed data-driven algorithm evaluates candidate controllers via scoring criteria that leverage system observability to isolate the effects of state history, enabling both detection of destabilizing controllers and accurate system identification. We present two algorithmic variants with dimension-free, finite-time guarantees, where each identifies the most suitable controller in $\mathcal{O}(N\log N)$ steps, while simultaneously achieving finite $L_{2}$-gain with respect to system disturbances.
00footnotetext: This work was supported by ONR grant #N00014-23-1-2299 and AFOSR MURI grant #FA9550-25-1-0375.00footnotetext: Corresponding author: Haoyuan Sun (haoyuans [at] mit [dot] edu).
@@ 1 Introduction
This paper presents a supervisory switching framework that guarantees finite-time adaptation for unknown linear systems, leveraging online algorithm to establish non-asymptotic analysis for this classical problem. In many applications involving complex systems, such as power systems [16], autonomous vehicles [2], and epidemic control [5], it is impractical for a single controller design to achieve acceptable performance across all operating conditions of the unknown plant. To address this challenge, switching control employs a collection of candidate controllers alongside a switching policy that selects the controller best suited to the current system behavior [14]. This approach has important applications in many established control paradigms, including gain scheduling [25, 23] and linear parameter-varying control [33, 17].
@@
To develop a non-asymptotic analysis of supervisory switching control, we revisit this classical problem through the lens of modern learning techniques. A natural starting point is the non-asymptotic analysis of system identification, which estimates the parameters of an unknown dynamical system from a single trajectory of input-output data. Conceptually, estimator-based supervisors select the model that minimizes the error between predicted and observed outputs, akin to the least-squares objectives underpinning system identification. Recent advances in non-asymptotic system identification have extended classical asymptotic results [15] to quantify the sample complexity of learning accurate system estimates from data [28, 34, 13]. This has led to numerous methods for the non-asymptotic identification of partially observed linear systems, typically by estimating the system’s response matrix to an exploratory signal and reconstructing the system parameters via the Ho-Kalman algorithm [21, 26, 24, 4]. However, these methods cannot be directly applied to our supervisory switching control paradigm because they presuppose system stability. This assumption is frequently violated in switching systems, where not all candidate controllers stabilize the underlying plant. Therefore, standard system identification approach cannot ensure safe exploration, as the potentially unstable controllers may cause the system to diverge before sufficient data is collected.
@@
In this work, we propose a data-driven analysis of supervisory switching control that establishes non-asymptotic guarantees for adapting to partially-observed linear systems with a collection of $N$ candidate controllers. Drawing on techniques from the machine learning and multi-armed bandit literature, our approach (Algorithm 1) evaluates each candidate controller over fixed switching intervals. Specifically, our evaluation criteria leverage system observability to approximate state trajectories from output data, effectively isolating the influences of earlier destabilizing controls. This eliminates the need for the stability or fully-observed state assumptions mandated by previous works.
@@ 3 Algorithm Design and Evaluation Criteria
2:Define $\bar{s_{i}}(q)=\frac{1}{q}\sum_{\ell=1}^{q}s_{i}[\ell]$ be the average score.
3:Let $Q_{i}(\ell)$ be the number of applications of controller $i$ through episode $\ell$.
@@
  • Theoretical analysis of online learning with states — such as reinforcement learning (RL) — typically rely on one of two assumptions to bound state trajectories during exploration. Episodic RL [9] assumes the system is periodically reset to a default initial state distribution, preventing compounding errors across episodes. Conversely, non-episodic online RL analyzes continuous operation without resets [27, 20, 30], but relies on restrictive structural assumptions, such as open-loop system stability, to prevent the state from diverging. In contrast, our setting requires exploring potentially destabilizing controllers without resets. This means that the systems states may diverge, thus violating the core assumptions of standard RL analysis.
@@ 3.1 Criterion 1: Instability Detection
For convenience, denote the current episode’s outputs as $y_{1},\dots,y_{\tau}$ and let $j$ index the active controller. We write out the following input-output relation for the first $\nu$ steps:
@@
We first consider the case where the active controller matches the true system ($i_{\star}=j$). In this case, the closed-loop dynamics are stable, and the Toeplitz matrices are bounded in norm by $\left\lVert C_{[j,j]},A_{[j,j]},B_{[j,j]}\right\rVert_{\mathcal{H}_{\infty}}$ (see [31]). Under Assumption 1, the closed-loop system is $\nu$-strictly observable, which means $\left\lVert\mathcal{O}_{[j,j]}^{\dagger}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon_{c}$. Consequently, the aggregated disturbance term $\mathcal{O}_{[j,j]}^{\dagger}\xi$ is sub-Gaussian. By standard concentration inequalities for sub-Gaussian random variables, the estimate
@@
Next, when the active controller is destabilizing, Assumption 2 ensures the closed-loop system possesses an explosive mode in the direction of the leading eigenvector of $A_{[i_{\star},j]}$. Since the system is persistently excited by Gaussian process noise, this explosive mode is activated with non-zero probability. In this regime, the residual $y_{t}-\hat{y}_{t}$ can be decomposed as the sum of estimation error plus accumulated disturbances, where the first part arises because $\hat{x}_{1}$ does not approximate the true initial state $x_{1}$ when $i_{\star}\neq j$. Critically, the estimation error is determined entirely by the first $\nu$ observations, while the system disturbances are accumulated at every step of the trajectory. As a result, the accumulated disturbance contains a component driven by process noise realized after the $\nu$-th step, which is independent of the estimation error. Hence, with non-zero probability, these independent components will not cancel, and contributions from the explosive mode will dominate the residual for sufficiently large $\tau$. Thus, the residual $y_{\tau}-\hat{y}_{\tau}$ will often exceed the threshold $\Theta$ when the active controller is destabilizing.
@@ 3.2 Criterion 2: System Identification
3:Compute a predicted trajectory
@@ 4 Main Result
With the guarantees of the two evaluation criteria established, we now analyze the performance of our supervisory switching control algorithm. To balance exploration and exploitation, we draw inspiration from the Upper Confidence Bound (UCB) algorithm [3]. At each episode, we select the controller by maximizing its empirical average score plus an exploration bonus. The magnitude of this bonus encourages exploration of controllers with limited interactions, and the parameter $a_{\ell}$ controls the exploration strength.
@@
Let $\mu^{+}$ denote a lower bound on the expected score of the matching controller ($j=i_{\star}$), and let $\mu^{-}$ denote an upper bound on the expected score of any non-matching controller ($j\neq i_{\star}$). Recall that $Q_{i}(\ell)$ denotes the number of times controller $i$ has been applied through episode $\ell$. Similar to standard UCB analysis, if Algorithm 1 selects a non-matching controller $j\neq i_{\star}$ at episode $\ell$, then at least one of the following three conditions must be met:
@@
  • The exploration bonus is too big: $\exists j\neq i_{\star}$, s.t. $2\sqrt{\frac{a_{\ell}}{Q_{j}(\ell)}}>\mu^{+}-\mu^{-}$.
@@
The first part of this theorem establishes a sample complexity bound, stating that our algorithm requires $\mathcal{O}(N\log N)$ episodes to identify the matching controller with high probability. This represents a significant improvement over our prior work [29] that required $\mathcal{O}(\exp(N))$ steps.
@@
In contrast, our Theorem 3 bounds the undiscounted case ($\lambda=0$). This regime is not covered by classical analysis, as the traditional framework requires a dwell time that scales inversely with $\lambda$. Furthermore, we can provide quantitative bounds for the constants $C_{0},C_{1}$ in terms of the problem parameters, whereas classical analysis is asymptotic and leaves the explicit values of $\overline{C}_{0}$ and $\overline{C}_{1}$ unspecified.
@@ 5 Conclusion
In this work, we presented a supervisory switching control framework that guarantees finite-time adaptation for unknown partially-observed linear systems with a collection of candidate controllers. Our proposed data-driven method achieves a dimension-free sample complexity of $\mathcal{O}(N\log N)$ steps to identify the matching controller while maintaining finite $L_{2}$-gain, even when some of the candidate controllers risk destabilizing the underlying plant.
@@ Appendix A Mathematical Preliminaries
Let $\{D_{k}\}_{k=1}^{\infty}$ be a martingale difference sequence adapted to a filtration $\{\mathcal{F}_{k}\}_{k=1}^{\infty}$. If $D_{k}\in[0,1]$ almost surely for all $k$, then for any $t>0$,
@@
Finally, we introduce the following technical lemma to manipulate inequalities of the form $b\geq a\log b$, which shows up when we bound the episode length $\tau$.
@@ Appendix B Proof of Proposition 1
3:Compute a predicted trajectory
@@
To start with the proof, we write out the following input-output relation for the first $\nu$ steps:
@@
$\hat{x}_{1}=\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}y_{\nu}\\ \vdots\\ y_{2}\\ y_{1}\end{bmatrix}=x_{1}+\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}\mathcal{T}^{(\nu-1)}_{[j,j]}\\ \mathbf{0}_{1\times(\nu-1)}\end{bmatrix}\begin{bmatrix}w_{\nu-1}\\ \vdots\\ w_{2}\\ w_{1}\end{bmatrix}+\mathcal{O}_{[j,j]}^{\dagger}\begin{bmatrix}\mathcal{T}^{(\nu-1)}_{[j,j]}\\ \mathbf{0}_{1\times(\nu-1)}\end{bmatrix}\begin{bmatrix}B_{[j,j]}u_{\nu-1}\\ \vdots\\ B_{[j,j]}u_{2}\\ B_{[j,j]}u_{1}\end{bmatrix}+\mathcal{O}_{[i_{\star},j]}^{\dagger}\begin{bmatrix}\eta_{\nu}\\ \vdots\\ \eta_{2}\\ \eta_{1}\end{bmatrix}$.
Under Assumption 1, the closed-loop system is $\nu$-strictly observable, and we have $\left\lVert\mathcal{O}_{[i_{\star},j]}^{\dagger}\right\rVert_{\mathrm{op}}\leq 1/\varepsilon_{c}$. And since the matching controller is stabilizing, the Toeplitz matrices are bounded by $\left\lVert C_{[j,j]},A_{[j,j]},B_{[j,j]}\right\rVert_{\mathcal{H}_{\infty}}$. It follows that $x_{1}-\hat{x}_{1}$ is
@@
Now, we can let $(u_{i_{\star}},v_{i_{\star}})$ be the leading singular vectors of $C_{[i_{\star},j]}A_{[i_{\star},j]}^{\tau-\nu-2}$ corresponding to the top singular value $\geq\varepsilon_{c}\varepsilon_{a}(1+\varepsilon_{a})^{\tau-2\nu-2}$. Since $w_{\nu+1}$ is Gaussian, we know that from standard normal CDF, $\Pr(|v_{i_{\star}}^{\top}w_{\nu+1}|>\sigma_{w}/4)>4/5$. Then we have that
@@
an unstable $A_{[i_{\star},j]}$ results in $\left\lVert y_{\tau}-\hat{y}_{\tau}\right\rVert\leq\sqrt{2d_{y}\Theta\log(2d_{y}/\delta)}$ with probability $\geq 2/5$. So, this criteria correctly detect an unstable $A_{[i_{\star},j]}$ with probability at least 2/5.
@@
an unstable $A_{[i_{\star},j]}$ results in $u_{i_{\star}}^{\top}(y_{\tau}-\hat{y}_{\tau})|\leq\sqrt{2d_{y}\Theta\log(2d_{y}/\delta)}$ with probability $\geq 2/5$. So, this criteria correctly detect an unstable $A_{[i_{\star},j]}$ with probability at least 2/5.
@@ Appendix C Proof of Proposition 2
Then, this criteria considers the active controller’s associated model as matching the true system if $\Delta_{u_{k},v_{k}}$ is small along every critical direction. And this criteria declares non-matching if any one of the $\Delta_{u_{k},v_{k}}$ is large.
@@ Appendix D Proof of Theorem 3
Given Proposition 1 and 2, we can instantiate the following guarantees for our evaluation metric: